Binocular TCS Testing with BP4: 19 October 2010 UT

Observer: JHill (LBTO) CBiddick, MDelapena (Tucson)
Telescope Operator: SAllanson, GBechetti (LBTO)
Telescope Support: JLittle (LBTO)
Instrument Support: JMorris (LBTO)

Summary

A fun night of binocular TCS testing with considerable progress. PCS is now sending optics offsets to the PSFs with the correct sign. Binocular pointing models seem to be working in a consistent way.

We also verified that RFBG is functional and working OK with the adaptive lookup table for astigmatism turned on. We did not spend any time testing the improved IDL WFS code.

Had no problems with Dave's binocular LBTtools. Dave tested it fully this afternoon.

We closed at 9:11 UT for high humidity.

Preparations

Chris has patched PSF to reverse the Ry M1M2 and call it Tilt for now and evermore.

Details

Open

Opened in twilight to a partly cloudy, but clearing sky.

02:14 Preset failed - rotator not ready

02:17 RIGHT - newpoint ACT0289 TelM=ACQUIRE gs=0 PAM- BC=1
LEFT - newpoint ACT0289 TelM=ACQUIRE gs=0 PAM- got a star

Tue Oct 19 02:17:44.215 2010  Preset alert for right IRTC mode ACQUIRE: dx.terc: [adjustCurrFocalStation.moveMirrorAbs.sendString]
Tue Oct 19 02:17:44.235 2010  Preset command failed for right IRTC mode ACQUIRE

The above error happened a couple times. Did we ever understand what it means?

ptspiral -s 0 sends values but doesn't move telescope mount, because we are only sending one side.

GCS acquisition windows are not labelled left/right.

SX Tip -15 SX Tilt -5

02:47 Binocular Active Preset gets to the acquisition images, but both sides fail at the start of guiding step (which is not unexpected at this point).

Right Side Monocular Tests with stars

Chris and Michele are keeping track of the tests. I'm just writing down what I did.

02:49 Set Monocular Mode on Right (Az dropped to vel mode - why?)

02:50 newpoint ACT0289 TelM=ACQUIRE gs=0 PAM- - failed for velmode

02:55 newpoint ACT0289 TelM=ACQUIRE gs=0 PAM-

ptspiral -d 1

ptspiral -s 1 -p 14 -d 4 -m 12 close at IE=-90 CA=-21 spiral from there -90 -7

03:15 why did the DX focus change? It seemed to shift by 0.3 mm without explanation.

DX IE=-67 CA+17

03:24 newpoint ACT0289 TelM=ACTIVE gs=13 PA=INDEF=18.08 - succeeded

03:27 focus 0 (Act Opt Offload is off)

03:31 offset 0 300 ABS RADEC

03:32 offset 0 0 ABS RADEC

03:34 newpoint ACT0289 TelM=ACTIVE gs=16 PA=90 (GS16 was too bright)

03:38 newpoint ACT0289 TelM=ACTIVE gs=17 PA=90 Roffset=5 Doffset=0 -succeeded

03:41 newpoint ACT0289 TelM=ACTIVE gs=13 PA=0 Roffset=5 Doffset=0

seeing is awful

03:45 newpoint ACT0289 TelM=ACTIVE gs=13 PA=0

03:47 offset 0 5 REL RADEC

03:48 offset 0 -10 REL RADEC

03:49 offset 0 -5 REL RADEC

03:50 offset 5 0 REL RADEC

03:51 offset -10 10 REL RADEC

03:56 newpoint ACT0289 TelM=ACTIVE gs=13 PA=0 XHot=1 YHot=1

03:58 newpoint ACT0289 TelM=ACTIVE gs=13 PA=0 XHot=3 YHot=3

Binocular Tests

04:02 Authorize IRTC on left (and right) side

04:05 RIGHT newpoint ACT0289 TelM=ACQUIRE gs=0 PAM- BC=1
LEFT newpoint ACT0289 TelM=ACQUIRE gs=0 PAM-
Got stars on both sides!

*** Pointing Model Terms *** Right
IA      -393.00000
IE       -67.00000
NPAE     -12.66810
CA        17.00000
AW        -6.65080
AN        22.89390
TF        26.40080
HACA2     -4.10620

*** Pointing Model Terms *** Left
IA      -264.00350
IE       -61.54510
NPAE       6.04950
AW        -3.43410
AN        22.53590
CA       -21.37560
TX        13.15460

SX Tip -11 Tilt -5

Repeat preset after changing IE CA

Mount moves if you make the same delta changes to IE CA on both sides

04:31 RIGHT newpoint ACT0289 TelM=ACTIVE gs=13 PA=0 BC=1
LEFT newpoint ACT0289 TelM=ACTIVE gs=13 PA=0
Acquisition image is being taken and displayed before telescope is in place.

04:33 Repeat the same preset with everything already there. Still the stars are not in the acquisition image.

04:42 RIGHT newpoint ACT0289 TelM=TRACK gs=0 PA=0 BC=1
LEFT newpoint ACT0289 TelM=TRACK gs=0 PA=0 - star appears on IRTC
Same result the Acq image is taken a second (modulo GCS display time) before the telescope settles on target.
Guiding almost works as the star was only a couple seconds outside the acq image

Oct 19 04:44:51.147: MCSPU: AZ/EL ON_SOURCE just became TRUE. AZ= 254.58911 EL=  66.56748
Oct 19 04:44:51.272: RFBG RTracker: Just changed to Rotator ON_SOURCE
Oct 19 04:44:51.337: LFBG RTracker: Just changed to Rotator ON_SOURCE

Oct 19 04:45:05 tcs4 LBT_right_GCS: full CCD readout requested, now waiting for image to arrive.
Oct 19 04:45:08 tcs4 LBT_left_GCS: full CCD readout requested, now waiting for image to arrive.

Oct 19 04:45:10 tcs4 LBT_right_GCS: acquisition image saved as /Repository/GCSfiles/dailyimages/acquire000014.fits
Oct 19 04:45:14 tcs4 LBT_left_GCS: acquisition image saved as /Repository/GCSfiles/dailyimages/acquire000015.fits

05:00 newpoint ACT0289 TelM=ACQUIRE gs=13 PA=0 BC=1 newpoint ACT0289 TelM=ACQUIRE gs=13 PA=0

05:16 Tip -14 Tilt -20

Michele patched PCS to have PCSutilities use side 2=both.

05:28 RIGHT newpoint ACT0289 TelM=ACTIVE gs=13 PA=0 BC=1
LEFT newpoint ACT0289 TelM=ACTIVE gs=13 PA=0

Oct 19 05:27:07.421: LFBG RTracker: Just changed to Rotator NOT ON_SOURCE !
Oct 19 05:27:07.436: MCSPU: AZ/EL ON_SOURCE just became FALSE AZ= 263.50409 EL=  57.81192
Oct 19 05:27:07.521: RFBG RTracker: Just changed to Rotator NOT ON_SOURCE !
Oct 19 05:27:07.822: RFBG RTracker: A slew appears to have started spontaneously while in TRACKING state
Oct 19 05:27:07.822: RFBG RTracker: State changed from TRACKING --> SLEW_TO_TRACK
Oct 19 05:27:07.874: LFBG RTracker: A slew appears to have started spontaneously while in TRACKING state
Oct 19 05:27:07.874: LFBG RTracker: State changed from TRACKING --> SLEW_TO_TRACK
Oct 19 05:27:07.973: RFBG RTracker: slew done.

Oct 19 05:27:07.973: RFBG RTracker: Slew done. state-->TRACKING, pos'n= 417847.6 v= -52.54
Oct 19 05:27:07.974: RFBG RTracker: State changed from SLEW_TO_TRACK --> TRACKING
Oct 19 05:27:08.025: LFBG RTracker: slew done.

Oct 19 05:27:08.025: LFBG RTracker: Slew done. state-->TRACKING, pos'n=1085028.7 v= -71.35
Oct 19 05:27:08.025: LFBG RTracker: State changed from SLEW_TO_TRACK --> TRACKING
Oct 19 05:27:08.712: AZ Tracker :slew done. position =  263.521deg. vel=     0.001 deg/sec

Oct 19 05:27:08.712: AZ Tracker :Track Poly: t0=4794182862.677 t1=4794182862.727 n=2 seq=1338102  263:31:10.3     10.37    -0.00
Oct 19 05:27:08.712: AZ Tracker :Slew done. state-->TRACKING, pos'n= 948673.9 v=   5.13
Oct 19 05:27:08.712: AZ Tracker :State changed from SLEW_TO_TRACK --> TRACKING
Oct 19 05:27:09.031: RFBG RTracker: Just changed to Rotator ON_SOURCE
Oct 19 05:27:09.082: LFBG RTracker: Just changed to Rotator ON_SOURCE
Oct 19 05:27:09.776: MCSPU: AZ/EL ON_SOURCE just became TRUE. AZ= 263.52236 EL=  57.79897

Oct 19 05:27:02 tcs4 LBT_right_GCS: full CCD readout requested, now waiting for image to arrive.
Oct 19 05:27:04 tcs4 LBT_left_GCS: full CCD readout requested, now waiting for image to arrive.

But GCSR hung here. ?

05:34 Stop and restart GCSR

05:40 Lower the PCS to optics threshold in lbt.conf from infinity to 0.25 arcsec. TipThreshold TiltThreshold (also have signs available)

*** Pointing Model Terms ***
LEFT
IA      -264.00350
IE       -56.00000
NPAE       6.04950
AW        -3.43410
AN        22.53590
CA       -47.00000
TX        13.15460
RIGHT
IA      -393.00000
IE       -74.00000
NPAE     -12.66810
CA       -10.00000
AW        -6.65080
AN        22.89390
TF        26.40080
HACA2     -4.10620

05:42 Stop/kill and restart PCS - takes minutes to connect to telemetry

RIGHT newpoint ACT0289 TelM=ACQUIRE gs=13 PA=0 BC=1
LEFT newpoint ACT0289 TelM=ACQUIRE gs=13 PA=0

Had to resend the RIGHT preset a couple times to avoid chase not idle due to optics updates.

05:56 Make pointing models on both sides IRTC_RFBG.20100527.ptmod

Modified ptspiral to do 2=both. It works OK, but it extracts the original IE/CA from left pointing model only. So they have to be the same for now/

06:03 RIGHT newpoint ACT0289 TelM=ACQUIRE gs=13 PA=0 BC=1
LEFT newpoint ACT0289 TelM=ACQUIRE gs=13 PA=0 - succeeded

Bad news: both telescopes have a 45 arcsec Ry correction at the start.

Good news: leaking works, as this common offset slowly bleeds away with each optics update.

06:17 reset both sides to the default pointing (not where my ptspiral left it)

ptspiral 7 arcsec pitch both This makes +-7 Tip Tilts appear on SX optics while running, but not on DX optics. We don't know why that is.

06:30 ptspiral -s 2 -p 12 -m 12 -d 4

Found a star on the DX guider at IE=-110 CA=-109, but nothing on IRTC, so it is the incorrect guide star.

06:43 Stop and restart IIF for a patch from Chris related to acquisition timing. register both sides in IRAF

06:46 RIGHT newpoint ACT0295 TelM=ACQUIRE gs=0 PAM- BC=1
LEFT newpoint ACT0295 TelM=ACQUIRE gs=0 PAM-
failed for several reasons

Goofy optics pointings were issued during the slew. 10^4

06:49 RIGHT newpoint ACT0295 TelM=ACQUIRE gs=0 PAM- BC=1
LEFT newpoint ACT0295 TelM=ACQUIRE gs=0 PAM-
- succeeded

We got a star on the left side at IE=-110 CA=-109 (maybe the wrong star)

06:52 reset both sides to the default pointing (not where my ptspiral left it)

06:54 ptspiral -s 2 -p 12 -m 12 -d 4

Found star on DX at IE=-96 CA=-77 with an offset of -12 on ??.

07:05 offset 0 0

Found DX at IE=-130 CA=-65

ptspiral -s 0 - can't find star

07:33 M5_0042 R=1.19 - failed, failed again both rotators faulted.

07:37 Restart MCSPU to clear the rotators.

Shell RIF'd (because we restarted the MCSPU)

Added Tilt = -15 M1M2 to DX Globals

07:58 Preset to ACT0305 which was a typo

08:00 Preset to ACT0325

Star is slowly moving down as Tip bleeds off optics, so sign of the Tip offset is wrong.

offset 0 5

08:11 Change the signs of both tip and tilt in lbt.conf
*** Pointing Model Terms ***
IA      -393.00000
IE       -45.00000
NPAE     -12.66810
CA        25.00000
AW        -6.65080
AN        22.89390
TF        26.40080
HACA2     -4.10620
IA      -393.00000
IE       -45.00000
NPAE     -12.66810
CA        25.00000
AW        -6.65080
AN        22.89390
TF        26.40080
HACA2     -4.10620

08:12 Stop and restart PCS

08:14 Preset to ACT0325 rotators drop to velocity mode

Az and El doing crazy stuff 08:19-08:20

Stop and restart PCS again

08:25 Preset to ACT0325

Probably flipping the signs also affected the PCS feedback loops as well as the output.

The shell Rifs whenever we stop MCSPU.

08:48 Start the patched PCS with the feedback corrected. Must revert SX pointing model and IE CA when you restart PCS.

08:52 RIGHT newpoint ACT0325 TelM=ACQUIRE gs=0 PAM- BC=1
LEFT newpoint ACT0325 TelM=ACQUIRE gs=0 PAM-

Found DX at IE=-70 CA=+25

09:08 RIGHT newpoint ACT0325 TelM=ACTIVE gs=1 BC=1
LEFT newpoint ACT0325 TelM=ACTIVE gs=1

Didn't find the star, but the timing looked better.

Good IE=-50 CA=+25 for the Right IRTC pointing model on both sides.

Close

09:11 Close for humidity at 94.6 and rising.

Revert TCS back to AO5

-- JohnHill - 19 Oct 2010
Topic revision: r2 - 19 Oct 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback